OpenStudy (anonymous):

Perform the operations and simplify when possible. −5(x − y) + 7(x − y) I keep getting 2x-12y, but that is incorrect

OpenStudy (anonymous):

-5x+5y+7x-7y=2x-2y=2(x-y) :)

OpenStudy (anonymous):

2x-2y